# A primer on inference and prediction with epidemic renewal models and sequential Monte Carlo

### v1

@@ 1 Introduction
Modern epidemiology relies on statistical models to track and project the spread of infectious diseases, inform public health policymaking, and estimate disease burden 1. A commonly used model for these purposes is the semi-mechanistic renewal model, which relates past incidence to current incidence through a simple renewal process 2. While this model is most commonly used for estimating the instantaneous reproduction number $R_{t}$ 3, 4, 5, it can also be used for generating projections 6, 7, 4, 8, estimating elimination probabilities 9, 10, 11, and modelling the effect of non-pharmaceutical interventions 12, 13, for example. Renewal models are frequently modified to account for various biases and limitations in epidemiological data 14, 15, 16, 17. These modifications often necessitate bespoke methods for fitting the model to data, which can be difficult to implement and become increasingly complex with additional modifications.
@@
Reviews of reproduction number estimation highlight key practical considerations to bear in mind when fitting the renewal model to epidemiological data. Gostic et al. (2020) 18 highlight generation interval misspecification, reporting delays, right-truncation due to observation delays, incomplete observation, and smoothing windows as particular concerns. Nash, Nouvellet, & Cori (2022) 14 identify 54 papers that make modifications to EpiEstim (arguably the most popular $R_{t}$ estimator) to account for such biases. They note delayed and missing reported case data, weekly reporting noise, alternative smoothing methods, imported cases, and temporal aggregation of reported data as common considerations. These adjustments are just as necessary for generating projections and other epidemiological analyses as they are for $R_{t}$ estimation.
@@
In this primer, we demonstrate the use of sequential Monte Carlo (SMC) methods 19, 20 as flexible and general-purpose solutions for fitting renewal models to data. These methods can handle multiple biases simultaneously and provide a unified framework for a range of tasks including reproduction number estimation and generating projections.
@@
SMC methods, often called particle filters, are a class of algorithms that are used to fit general hidden-state models to observed data. They use a collection of samples (called particles) to represent the posterior distribution of “hidden states” at each time step (e.g. the instantaneous reproduction number or daily infection incidence) given some observed data (e.g. reported cases). These particles are projected forward according to a state-space transition model (typically an epidemic model of some kind) and weighted by an observation model (typically representing a noisy observation process). Such methods are particularly powerful, as they only require simulations from the epidemic model, rather than evaluations of a likelihood, allowing for simulation-based inference 21. Accounting for the aforementioned delays and biases simply requires incorporating these in the state-space transition and observation models, allowing a wide range of models to be fit using a single framework.
@@
Despite being well suited to fitting epidemiological models and having a relatively straightforward and intuitive implementation, and calls for their further use 22, SMC methods are not widely used when fitting epidemic renewal models. They have seen greater use in fitting compartmental models 23, 24, 25, 25, 26, 27, 28, 29, while examples of their application to agent-based models 30, Hawkes processes 31, and renewal models 32, 33, 34 are more limited. This discrepancy likely arises from the Markovian nature of many compartmental models, a property which is leveraged by many popular SMC methods, whereas the renewal model is explicitly non-Markovian. We are aware of two reviews of SMC methods in epidemiology 35, 36, both of which are high quality and provide complementary information to this primer, though neither covers methods applicable to the renewal model. To aid comparison with these papers, a discussion of similarities and differences with this primer is provided online.
@@
Methods for fitting renewal models that feature comparable flexibility typically rely on Hamiltonian Monte Carlo (HMC)4, 37, 38 (as implemented in Stan 39), often coupled with an approximate Gaussian process state-space model (such as in EpiNow2 4, which is available as a plug-and-play software package). However, HMC-based methods do not support discrete state spaces or parameters, can be comparatively computationally intensive, and cannot be updated online as new data are observed. Additionally, Gaussian process approximations introduce significant mathematical complexity. Our goal is to enable a researcher to rapidly implement their own methods, ensuring they understand the process from start to finish. An alternative flexible approach is the Laplacian P-spline method of EpiLPS 40, though its fully Bayesian implementation is also gradient-based, again limiting it to continuous state spaces and parameters, and the mathematical complexity of this method is also high. Finally, approximate Bayesian computation methods could also be used to fit renewal models, although (as for SMC) the literature is largely focused on compartmental models 41, 42.
@@
We aim to familiarise the reader with both the renewal model and SMC methods. We provide simple implementations of SMC-type algorithms, including a bootstrap filter with fixed-lag resampling and a particle marginal Metropolis-Hastings (PMMH) algorithm, for fitting renewal models to data. Their use is demonstrated in several epidemiologically interesting scenarios using data from the COVID-19 pandemic in Aotearoa New Zealand. We also provide guidance on model evaluation and comparison. Our aim is not to provide a model for every scenario, but to equip the reader with the tools to construct and fit their own models.
@@
This paper contains the technical details needed to construct and fit epidemic renewal models with SMC methods. Some readers may find it helpful to consult the example code online while reading, which is available as downloadable notebooks from SMC and epidemic renewal models. This website also contains additional examples, tutorials, and longer explanations of the topics considered here.
@@ 2 Hidden-state models
A typical hidden-state model consists of time-indexed unobserved hidden states $X_{t}$, observed data $y_{t}$, and parameter(s) $\theta$. These may be scalar or vector-valued. The hidden states represent the underlying process of interest (e.g. the true infection incidence and reproduction number) from which the observed data (e.g. reported cases) are generated. The goal is to infer the value of the hidden states $X_{t}$ and/or parameters $\theta$ given the observed data. SMC methods, discussed later, are a class of algorithms for fitting general hidden-state models to observed data 19.
@@
The notation $X_{s:t}$ denotes the values of this quantity at all time steps from $s$ to $t$, inclusive. In epidemiological applications, this distribution is typically implied by an epidemic model describing how the epidemic evolves over time. Stochastic compartmental models are a popular choice here 35, 36, although other models (such as the renewal model) can be used.
@@
Other names for hidden-state models include Hidden Markov Models (HMMs), when the model is Markovian 43; partially observed Markov processes (POMPs), also for Markovian models 44; and latent variable models, where $X_{t}$ are referred to as latent variables 45. The renewal model we examine is not Markovian, as infection incidence today depends on a convolution of infection incidence across multiple previous time steps, thus it cannot be treated as an HMM or POMP. Augmenting past states into $X_{t}$, i.e., setting $X_{t}=(\tilde{X}_{t},\tilde{X}_{t-1},\ldots)$ can transform a non-Markovian model into a Markovian one, at the cost of an enlarged hidden-state space. We do not consider this further in this primer.
Many popular epidemic models can be framed as a hidden-state model. EpiFilter, which can be used to estimate the instantaneous reproduction number $R_{t}$ 5 or infer the elimination probability of an epidemic 9, is explicitly formulated this way. In this model, $R_{t}$ is assumed to follow a Gaussian random walk (the state-space transition model) and observed cases $C_{t}$ are assumed to follow the Poisson renewal model (the observation model). A less obvious example of a hidden-state model is that by Fraser (2007) 46, which shares EpiFilter’s observation model but assumes $R_{t}$ is piecewise constant on intervals of 10 days. Further examples include the various implementations of EpiNow2 4, and a model for temporally aggregated and underreported data by Ogi-Gittins et al. (2025)47. These models all rely on the renewal model and employ bespoke inference methods, such as a grid-based approximation to the Bayesian filtering and smoothing equations in EpiFilter, maximum likelihood estimation in Fraser (2007) 46, or HMC in EpiNow2 4.
@@ 3 The renewal model
While SMC methods can be used to fit various hidden-state models, and many epidemic models can be formulated as hidden-state models, we focus on the epidemic renewal model in this primer.
Denoting reported cases at time $t$ by $C_{t}$, the reproduction number by $R_{t}$, and the probability that a secondary case is reported $u$ days after the primary case by $\omega_{u}$ (often approximated by the probability mass function (PMF) of the serial interval), the renewal model relates reported cases today to previously reported cases through the renewal equation:
@@
If $C_{t}$ follows a Poisson distribution with mean $E[C_{t}]$, then we have the popular Poisson renewal model. However, other distributions can be used 48, often to account for overdispersion in transmission or “superspreading” 49. The serial interval is the time between the symptom onset of the primary case and the symptom onset of the secondary case, the PMF of which is represented by $\{\omega_{u}\}_{u=1}^{u_{max}}$. The term $C_{t-u}\omega_{u}$ in the summation represents the expected contribution to today’s reported cases by reported cases from $u$ time steps ago. This relationship is visualised in figure 2.
In a simple case, $R_{t}$ follows a Gaussian random walk (the state-space transition model), and daily reported cases have an expected value defined by equation (1) (the observation model). If a Poisson observation distribution is used, then this is the model defined by EpiFilter 5, which we also adopt in example model 1.
A more realistic model might assume that underlying infection incidence follows the renewal model: equation 1 is modified by replacing observed $C_{t}$ with unobserved $I_{t}$, and the serial interval distribution with the generation time distribution. Reported cases are then treated as noisy observations of infection incidence. In this setup, the renewal model now defines part of the state-space transition model, and the observation process is modelled separately. This is the approach adopted in example models 2 and 3. Crucially, SMC methods can be used in both cases, whereas EpiFilter is limited to the case where the renewal model defines the observation distribution.
@@
Figure 2: Diagram of the renewal model. In this example, the serial interval takes values $\omega_{u}=0.25,0.5,0.25$ for $u=1,2,3$ (a maximum serial interval of three days) and the reproduction number is assumed to take values $R_{t-1}=2$, $R_{t}=2$, and $R_{t+1}=1$. On the left, the expected number of secondary cases produced by a primary case who was reported at time $t\!-\!2$ is shown (0.5 cases on day $t\!-\!1$, a single case on day $t$, and 0.25 cases on day $t\!+\!1$), with their expected contribution to total cases at time $t$ highlighted. On the right, the expected total cases at time $t$ is shown as the sum of the expected cases produced by primary cases reported $u=1,2,\text{ and }3$ days ago, defining the renewal equation.
@@ 4.1 The bootstrap filter
The goal of the bootstrap filter 50 is to generate samples from the filtering and/or smoothing distributions, conditional on the observed data $y_{1:T}$ and parameter vector $\theta$. This is a similar goal to EpiFilter, for example, which targets the posterior distribution of $R_{t}$ given observed cases $C_{1:T}$ and parameter $\eta$ 5. The bootstrap filter is one of the simplest particle filtering methods, although more advanced methods exist and may be preferable in some settings 51. The similarity to EpiFilter, which produces analytical solutions, offers a convenient way to check a bootstrap filter implementation: fitting the same model to the same data allows outputs to be compared directly.
@@
The algorithm starts with an initial set of $N$ “particles” $\{x_{0}^{(i)}\}_{i=1}^{N}$ sampled from an arbitrary initial-state distribution $P(X_{0})$. These particles are projected forward by sampling from the state-space transition distribution $\tilde{x_{t}}^{(i)}\sim P(X_{t}|x_{0:t-1},\theta)$, representing equally likely one-step-ahead projections of the hidden states before any data are observed. The projected particles are then weighted by the observation model $w_{t}^{(i)}=P(y_{t}|\tilde{x_{t}}^{(i)},y_{1:t-1},\theta)$, quantifying how likely each projected particle is once the data at time step $t$ are observed. The particles are then resampled with replacement according to these weights, resulting in a new set of particles $\{x_{t}^{(i)}\}_{i=1}^{N}$ that represent samples from the posterior distribution at time step $t$. This process is described in figure 4. Repeating for $t=1,\ldots,T$ completes the algorithm, described formally in algorithm 1. Further intuition in the language of sequential importance sampling is provided online.
@@
Figure 4: A single step of the bootstrap filter with resampling. Particles are projected forward by sampling according to the state-space transition distribution, weighted by the observation distribution, and resampled to form a new set of particles. In this example, the orange and yellow trajectories receive zero weight, so do not feature in the right-hand panel. The remaining trajectories feature in the right-hand panel with relative frequency proportional to their weight.
@@
If only the particles $x_{t}^{(i)}$ at time step $t$ are resampled, then samples from the filtering distributions $P(X_{t}|y_{1:t},\theta)$ are returned (for each $t$). If all particle histories $x_{1:t}^{(i)}$ are resampled at each time step (as in figure 4), then samples from the smoothing distributions $P(X_{t}|y_{1:T},\theta)$ are returned. Intuitively, this is because resampled past particles are effectively re-weighted according to current data.
@@
When $T$ is large, resampling entire particle histories leads to particle degeneracy 20, where few unique particles remain at early time steps. This is visible in figure 4 as a decreasing number of unique particle trajectories after resampling, which results in poor approximations of the smoothing distribution. To mitigate particle degeneracy, we use fixed-lag resampling 51, resampling the past $L$ time steps $x_{t-L:t}^{(i)}$ rather than the full state histories, which yields an approximation to the smoothing distribution: $P(X_{t}|y_{1:t+L},\theta)$. If $X_{t}$ is conditionally independent of data more than $L$ steps in the future, given $y_{1:t+L-1}$, then this approximation is exact. In renewal models, $L$ should exceed the maximum serial interval, and, if reporting delays are present, $L$ should also exceed the maximum reporting delay.
@@
Algorithm 1 The fixed-lag bootstrap filter. Statements indexed by $i$ should be taken to mean for $i$ in $\{1,2,\ldots,N\}$. At time step $t$, $\tilde{x}_{i,t-L:t}$ is taken to mean the aggregation of particle history $x_{i,t-L:t-1}$ and newly projected particles $\tilde{x}_{i,t}$.
@@
4:     $\tilde{x}_{i,t}\sim P(X_{t}|x_{i,0:t-1},\theta)$triangleright Sample from the state-space transition distribution
5:     $w_{i,t}\leftarrow P(y_{t}|\tilde{x}_{i,t-L:t},y_{1:t-1},\theta)$triangleright Calculate the observation weights
6:     $x_{i,t-L:t}\sim\text{Multinomial}\left(\{\tilde{x}_{j,t-L:t}\}_{j=1}^{N},\{w_{j,t}\}_{j=1}^{N}\right)$triangleright Resample the particles
@@
8:Return: Particle values $x_{i,t}$ and observation weights $w_{i,t}$ for $i=1,\ldots,N$ and $t=1,\ldots T$.

### v2

@@ 1 Introduction
Modern epidemiology relies on statistical models to track and project the spread of infectious diseases, inform public health policymaking, and estimate disease burden 1. A commonly used model for these purposes is the semi-mechanistic renewal model, which relates past incidence to current incidence through a simple renewal process 2, 3, 4. While this model is most commonly used for estimating the instantaneous reproduction number $R_{t}$ 5, 6, 7, it can also be used for generating projections 8, 9, 6, 10, estimating elimination probabilities 11, 12, 13, and modelling the effect of non-pharmaceutical interventions 14, 15, for example. Renewal models are frequently modified to account for various biases and limitations in epidemiological data 16, 17, 18, 19. These modifications often necessitate bespoke methods for fitting the model to data, which can be difficult to implement and become increasingly complex with additional modifications.
@@
Reviews of reproduction number estimation highlight key practical considerations to bear in mind when fitting the renewal model to epidemiological data. Gostic et al. (2020) 20 highlight generation interval misspecification, reporting delays, right-truncation due to observation delays, incomplete observation, and smoothing windows as particular concerns. Nash, Nouvellet, & Cori (2022) 16 identify 54 papers that make modifications to EpiEstim (arguably the most popular $R_{t}$ estimator) to account for such biases. They note delayed and missing reported case data, weekly reporting noise, alternative smoothing methods, imported cases, and temporal aggregation of reported data as common considerations. These adjustments are just as necessary for generating projections and other epidemiological analyses as they are for $R_{t}$ estimation.
@@
In this primer, we present a pedagogical overview of how sequential Monte Carlo (SMC) methods 21, 22, which are well established in other contexts, can be used as flexible and general-purpose tools for fitting renewal models to epidemiological data. The overall approach can be viewed as a distillation of the methods employed by Watson et al. (2024) 23. We focus on how these methods can handle multiple biases simultaneously and provide a unified framework for a range of tasks including reproduction number estimation and generating projections.
@@
SMC methods, often called particle filters, are a class of algorithms that are used to fit general hidden-state models to observed data. They use a collection of samples (called particles) to represent the posterior distribution of “hidden states” at each time step (e.g. the instantaneous reproduction number or daily infection incidence) given some observed data (e.g. reported cases). These particles are projected forward according to a state-space transition model (typically an epidemic model of some kind) and weighted by an observation model (typically representing a noisy observation process). Such methods are particularly powerful, as they only require simulations from the epidemic model, rather than evaluations of a likelihood, allowing for simulation-based inference 24. Accounting for the aforementioned delays and biases simply requires incorporating these in the state-space transition and observation models, allowing a wide range of models to be fit using a single framework.
@@
Despite being well suited to fitting epidemiological models and having a relatively straightforward and intuitive implementation, and calls for their further use 25, SMC methods are not widely used when fitting epidemic renewal models. They have seen greater use in fitting compartmental models 26, 27, 28, 28, 29, 30, 31, 32, while examples of their application to agent-based models 33, Hawkes processes 34, and renewal models 35, 23, 36 are more limited. This discrepancy likely arises from the Markovian nature of many compartmental models, a property which is leveraged by many popular SMC methods, whereas the renewal model is explicitly non-Markovian. We are aware of two reviews of SMC methods in epidemiology 37, 38, both of which are high quality and provide complementary information to this primer, though neither covers methods applicable to the renewal model. To aid comparison with these papers, a discussion of similarities and differences with this primer is provided online.
@@
Methods for fitting renewal models that feature comparable flexibility typically rely on Hamiltonian Monte Carlo (HMC)6, 39, 40 (as implemented in Stan 41), often coupled with an approximate Gaussian process state-space model (such as in EpiNow2 6, which is available as a plug-and-play software package). However, HMC-based methods do not support discrete state spaces or parameters, can be comparatively computationally intensive, and cannot be updated online as new data are observed. Additionally, Gaussian process approximations introduce significant mathematical complexity. Our goal is to enable a researcher to rapidly implement their own methods, ensuring they understand the process from start to finish. An alternative flexible approach is the Laplacian P-spline method of EpiLPS 42, though its fully Bayesian implementation is also gradient-based, again limiting it to continuous state spaces and parameters, and the mathematical complexity of this method is also high. Finally, approximate Bayesian computation methods could also be used to fit renewal models, although (as for SMC) the literature is largely focused on compartmental models 43, 44.
@@
We aim to familiarise the reader with both the discrete-time renewal model and SMC methods. We provide simple implementations of SMC-type algorithms, including a bootstrap filter with fixed-lag resampling and a particle marginal Metropolis-Hastings (PMMH) algorithm, for fitting renewal models to data. Their use is demonstrated in several epidemiologically interesting scenarios using national data from the COVID-19 pandemic in Aotearoa New Zealand. We also provide guidance on model evaluation and comparison. Our aim is not to provide a model for every scenario, but to equip the reader with the tools to construct and fit their own models.
@@
This paper contains the technical details needed to construct and fit discrete-time epidemic renewal models with SMC methods, consolidating established methods into a single, accessible framework. Some readers may find it helpful to consult the example code online while reading, which is available as downloadable notebooks from SMC and epidemic renewal models. This website also contains additional examples, tutorials, and longer explanations of the topics considered here.
@@ 2 Hidden-state models
A typical hidden-state model consists of time-indexed unobserved hidden states $X_{t}$, observed data $y_{t}$, and parameter(s) $\theta$. These may be scalar or vector-valued. The hidden states represent the underlying process of interest (e.g. the true infection incidence and reproduction number) from which the observed data (e.g. reported cases) are generated. The goal is to infer the value of the hidden states $X_{t}$ and/or parameters $\theta$ given the observed data. SMC methods, discussed later, are a class of algorithms for fitting general hidden-state models to observed data 21.
@@
The notation $X_{s:t}$ denotes the values of this quantity at all time steps from $s$ to $t$, inclusive. In epidemiological applications, this distribution is typically implied by an epidemic model describing how the epidemic evolves over time. Stochastic compartmental models are a popular choice here 37, 38, although other models (such as the renewal model) can be used.
@@
Other names for hidden-state models include Hidden Markov Models (HMMs), when the model is Markovian 45; partially observed Markov processes (POMPs), also for Markovian models 46; and latent variable models, where $X_{t}$ are referred to as latent variables 47. The renewal model we examine is not Markovian, as infection incidence today depends on a convolution of infection incidence across multiple previous time steps, thus it cannot be treated as an HMM or POMP. Augmenting past states into $X_{t}$, i.e., setting $X_{t}=(\tilde{X}_{t},\tilde{X}_{t-1},\ldots)$ can transform a non-Markovian model into a Markovian one, at the cost of an enlarged hidden-state space. We do not consider this further in this primer.
Many popular epidemic models can be framed as a hidden-state model. EpiFilter, which can be used to estimate the instantaneous reproduction number $R_{t}$ 7 or infer the elimination probability of an epidemic 11, is explicitly formulated this way. In this model, $R_{t}$ is assumed to follow a Gaussian random walk (the state-space transition model) and observed cases $C_{t}$ are assumed to follow the Poisson renewal model (the observation model). A less obvious example of a hidden-state model is that by Fraser (2007) 48, which shares EpiFilter’s observation model but assumes $R_{t}$ is piecewise constant on intervals of 10 days. Further examples include the various implementations of EpiNow2 6, and a model for temporally aggregated and underreported data by Ogi-Gittins et al. (2025)49. These models all rely on the renewal model and employ bespoke inference methods, such as a grid-based approximation to the Bayesian filtering and smoothing equations in EpiFilter, maximum likelihood estimation in the model by Fraser et al. (2011) 48, or HMC in EpiNow2 6.
@@ 3 The renewal model
While SMC methods can be used to fit various hidden-state models, and many epidemic models can be formulated as hidden-state models, we focus on the discrete-time epidemic renewal model in this primer. The renewal model was introduced by Leonhard Euler in 1767 and popularised in the contemporary continuous-time formulation by Alfred Lotka in 1907, who used it to model age-structured population dynamics 50. This continuous-time formulation was first introduced into epidemiology alongside the susceptible-infected-recovered (SIR) model by Kermack and McKendrick51 and gained wider popularity during the 1970s 2. The modern discrete-time formulation was proposed by Fraser (2007)3 and subsequently popularised as a Bayesian estimator of $R_{t}$ in Cori et al. (2013) 52, 5. Since then, the discrete-time renewal model has become a popular semi-mechanistic model of disease transmission 4.
Denoting reported cases at time $t$ by $C_{t}$, the reproduction number by $R_{t}$, and the probability that a secondary case is reported $u$ time steps after the primary case by $\omega_{u}$ (often approximated by the probability mass function (PMF) of the serial interval), the renewal model relates reported cases at time $t$ to previously reported cases through the renewal equation:
@@
If $C_{t}$ follows a Poisson distribution with mean $E[C_{t}]$, then we have the popular Poisson renewal model. However, other distributions can be used 53, often to account for overdispersion in transmission or “superspreading” 54. The serial interval is the time between the symptom onset of the primary case and the symptom onset of the secondary case, the PMF of which is represented by $\{\omega_{u}\}_{u=1}^{u_{max}}$, where $u_{max}$ is the assumed maximum possible serial interval. The term $C_{t-u}\omega_{u}$ in the summation represents the expected contribution to current reported cases by reported cases from $u$ time steps ago. This relationship is visualised in Figure 2.
@@
Figure 2: Diagram of the renewal model. In this example, the serial interval takes values $\omega_{u}=0.25,0.5,0.25$ for $u=1,2,3$ (a maximum serial interval of three time steps) and the reproduction number is assumed to take values $R_{t-1}=2$, $R_{t}=2$, and $R_{t+1}=1$. On the left, the expected number of secondary cases produced by a primary case who was reported at time $t\!-\!2$ is shown (0.5 cases at time $t\!-\!1$, a single case at time $t$, and 0.25 cases at time $t\!+\!1$), with their expected contribution to total cases at time $t$ highlighted. On the right, the expected total cases at time $t$ is shown as the sum of the expected cases produced by primary cases reported $u=1,2,\text{ and }3$ time steps ago, defining the renewal equation.
In a simple example, $R_{t}$ follows a Gaussian random walk (the state-space transition model), and daily reported cases have an expected value defined by equation (1) (the observation model). If a Poisson observation distribution is used, then this is the model defined by EpiFilter 7, which we also adopt in example model 1.
A more realistic model might assume that underlying infection incidence follows the renewal model: equation 1 is modified by replacing observed $C_{t}$ with unobserved $I_{t}$, and the serial interval distribution with the generation time distribution. Reported cases are then treated as noisy observations of infection incidence. In this setup, the renewal model now defines part of the state-space transition model, and the observation process is modelled separately. Explicitly modelling infection-to-infection, instead of reported case-to-reported case, has the added benefit of allowing for potentially negative serial intervals, which would otherwise be difficult to account for, especially when fitting models to paired data with observed negative intervals. This is the approach adopted in example models 2 and 3. Crucially, SMC methods can be used in both cases, whereas EpiFilter is limited to the case where the renewal model defines the observation distribution.
The discrete-time nature of our model necessitates the specification of an appropriate time-step duration. The key constraint can be seen in Equation 1, which assumes that primary cases (or infections if we model the underlying infection incidence) cannot produce secondary cases (secondary infections) on the same day they were reported (infected). The discretisation should be fine enough that the probability of same-day transmission is negligible 55. For COVID-19, daily time steps were popular, as this coincided with the frequency of reported case data.
In some scenarios, data may be reported on a longer time scale than infections occur. For example, weekly reporting is common in influenza surveillance, while typical influenza generation times are measured in days 56. Published literature accounts for this using simulation-based methods 55, 56, expectation-maximisation 57, or Gaussian processes 6 to infer infection incidence at a finer time-scale before applying the renewal model. The SMC approach introduced in this primer does not assume data are observed on every time step, allowing an appropriate discretisation to be chosen to suit the pathogen being modelled, independently of how frequently data are reported. This is examined in a practical example in Section 6.3. We also demonstrate this on an influenza dataset online.
@@ 4.1 The bootstrap filter
The goal of the bootstrap filter 58 is to generate samples from the filtering and/or smoothing distributions, conditional on the observed data $y_{1:T}$ and parameter vector $\theta$. This is a similar goal to EpiFilter, for example, which targets the posterior distribution of $R_{t}$ given observed cases $C_{1:T}$ and parameter $\eta$ 7. The bootstrap filter is one of the simplest particle filtering methods, although more advanced methods exist and may be preferable in some settings 59. The similarity to EpiFilter, which produces analytical solutions, offers a convenient way to check a bootstrap filter implementation: fitting the same model to the same data allows outputs to be compared directly.
@@
The algorithm starts with an initial set of $N$ “particles” $\{x_{0}^{(i)}\}_{i=1}^{N}$ sampled from an arbitrary initial-state distribution $P(X_{0})$. These particles are projected forward by sampling from the state-space transition distribution $\tilde{x_{t}}^{(i)}\sim P(X_{t}|x_{0:t-1},\theta)$, representing equally likely one-step-ahead projections of the hidden states before any data are observed. The projected particles are then weighted by the observation model $w_{t}^{(i)}=P(y_{t}|\tilde{x_{t}}^{(i)},y_{1:t-1},\theta)$, quantifying how likely each projected particle is once the data at time step $t$ are observed. The particles are then resampled with replacement according to these weights, resulting in a new set of particles $\{x_{t}^{(i)}\}_{i=1}^{N}$ that represent equally-weighted samples from the posterior distribution at time step $t$. This process is described in Figure 4. Repeating for $t=1,\ldots,T$ completes the algorithm, described formally in Algorithm 1. Further intuition in the language of sequential importance sampling is provided online. Readers familiar with alternative SMC methods may expect all weights to be reset to $1/N$ following resampling. This is unnecessary as we resample at every step, thus we always treat the resampled particles as equally-weighted samples from the target posterior distribution and past weights do not feature in calculations at future time steps (that is, there is nothing to “reset”). Furthermore, the unnormalised weights $w_{t}^{(i)}$ are later used when estimating the likelihood in Algorithm 2, so it is helpful to store these as initially calculated.
@@
Figure 4: A single step of the bootstrap filter with resampling. Particles are projected forward by sampling according to the state-space transition distribution, weighted by the observation distribution, and resampled to form a new set of equally-weighted particles. In this example, the orange and yellow trajectories receive zero weight, so do not feature in the right-hand panel. The remaining trajectories feature in the right-hand panel with relative frequency proportional to their weight.
@@
If only the particles $x_{t}^{(i)}$ at time step $t$ are resampled, then equally-weighted samples from the filtering distributions $P(X_{t}|y_{1:t},\theta)$ are returned (for each $t$). If all particle histories $x_{1:t}^{(i)}$ are resampled at each time step (as in Figure 4), then equally-weighted samples from the smoothing distributions $P(X_{t}|y_{1:T},\theta)$ are returned. Intuitively, this is because resampling effectively re-weights past particles according to current data.
@@
When $T$ is large, resampling entire particle histories leads to particle degeneracy 22, where few unique particles remain at early time steps. This is visible in Figure 4 as a decreasing number of unique particle trajectories after resampling, which results in poor approximations of the smoothing distribution. To mitigate particle degeneracy, we use fixed-lag resampling 59, resampling the past $L$ time steps $x_{t-L:t}^{(i)}$ rather than the full state histories, which yields an approximation to the smoothing distribution: $P(X_{t}|y_{1:t+L},\theta)$. If $X_{t}$ is conditionally independent of data more than $L$ steps in the future, given $y_{1:t+L-1}$, then this approximation is exact. In renewal models, $L$ should exceed the maximum serial interval $u_{max}$, and, if reporting delays are present, $L$ should also exceed the maximum reporting delay.
@@
Algorithm 1 The fixed-lag bootstrap filter. Statements indexed by $i$ should be taken to mean for $i$ in $\{1,2,\ldots,N\}$. At time step $t$, $\tilde{x}_{t-L:t}^{(i)}$ is taken to mean the aggregation of particle history $x_{t-L:t-1}^{(i)}$ and newly projected particles $\tilde{x}_{t}^{(i)}$.
@@
4:     $\tilde{x}_{t}^{(i)}\sim P(X_{t}|x_{0:t-1}^{(i)},\theta)$triangleright Sample from the state-space transition distribution
5:     $w_{t}^{(i)}\leftarrow P(y_{t}|\tilde{x}_{t-L:t}^{(i)},y_{1:t-1},\theta)$triangleright Calculate the observation weights
6:     $x_{t-L:t}^{(i)}\sim\text{Multinomial}\left(\{\tilde{x}_{t-L:t}^{(j)}\}_{j=1}^{N},\left\{\frac{w_{t}^{(j)}}{\sum_{k=1}^{N}w_{t}^{(k)}}\right\}_{j=1}^{N}\right)$triangleright Resample the particles
@@
8:Return: Particle values $x_{t}^{(i)}$ and observation weights $w_{t}^{(i)}$ for $i=1,\ldots,N$ and $t=1,\ldots T$.